OpenStudy (anonymous):

Determine the exact value for cos^-1(cos(7π/6)). ...& Also... Determine the value of cos(-7π/4).

OpenStudy (anonymous):

Well \[\Large \cos (\cos^{-1} (x))\]

OpenStudy (anonymous):

=\[\Large x\]

OpenStudy (anonymous):

\[\Large \cos(-x)=\cos(x)=\cos(-\frac{7\pi}{4})=\cos(\frac{7\pi}{4})\]

OpenStudy (anonymous):

You can proceed from there right?

OpenStudy (kainui):

Since cos^-1() is the inverse function of cos(), you're just undoing it. It's like squaring the square root function, you get back what's inside. To find the value of cos(-7π/4) you should realize that cosine is basically looking for the x-value of rotation on a circle and that a circle has 2π radians in all. since they say -7π/4, that means you rotate around clockwise because of the negative sign for 7 quarters of pi.